# A Population Analysis of 20 Exoplanets Observed from the Optical to the Near-infrared Wavelengths with HST: Evidence for Widespread Stellar Contamination

### v2

@@ 5.3 SAD metric results
As discussed in Section 5.2, the Bayesian approach favours precise data, giving less weight to data points with larger uncertainties. Attempting a retrieval on the spectral bands covered by STIS data only is not advisable, as the inclusion of WFC3 data is necessary to anchor the spectra to the continuum level. However, we still want to assess how well the retrieval captures potential stellar activity signals in the STIS data sets. Hence, we developed the SAD, a metric that assesses the goodness-of-fit in the optical wavelengths that acts independently of Bayesian statistics. A description of the metric is given in Section 4.2 while the results are reported in Fig. 5 and 6 for Case 1 and Case 2 respectively, in ascending SAD order. The SAD results for GJ 3470 b Case 3 and HD 209458 b Case 3 and 4 are displayed in Fig. 7.
From the SAD Case 1 results we can infer that WASP-127 b is the planet for which the stellar contamination model does the poorest job at fitting the optical data (likely due to the large spread of the datapoints), while the strong optical slope seen in WASP-6 b’s STIS data is substantially better described by the activity model compared to the base model. The SAD value obtained for WASP-127 b is 0.8440.315 compared to 2.9741.652 for WASP-6 b, a difference exceeding three times the value, although we note that the uncertainty on WASP-6 b is unavoidably large due to the comparatively large error bars on the data points themselves. For WASP-127 b, the stellar model is better at capturing the small-scale spectral modulations (for instance between 0.4 and 0.5 m), however despite this, the mean distance between the best-fit model and the observed data points is ultimately minimised with the base model, although arguably neither model fits the observation particularly well. The SAD is designed to support the model that, on average, passes closer to the highest number of data points. An inspection of the WASP-127 b spectrum confirms this: even though the base model is able to fit seven data points within their 1 error bars, compared to nine data points for the stellar model, the average transit depth distances from the base model are smaller than the average distances from the stellar model. This is how the base model prevails.
Between WASP-127 b and WASP-6 b, a diverse range of SADs is observed, reflecting the diversity of spectral shapes and the models’ ability to fit the data. Six planets, HD 209458 b, WASP-69 b, HAT-P-11 b, HAT-P-12 b, WASP-31 b, WASP-29 b, have a SAD around 1 (0.990SAD1.10), which suggests a negligble/weak preference for either model. WASP-127 b and WASP-39 b mostly favour the base model (SAD 0.990), while GJ 3470 b, GJ-3470 b, HAT-P-1 b, WASP-101 b, WASP-80 b, WASP-17 b, HAT-P-18 b, WASP-121 b, HAT-P-26 b, WASP-74 b, WASP-52 b and WASP-6 b are all best described by an active star model (SAD higher than 1.10).
Focusing on the results from Case 2 and employing the same criteria as Case 1, we identify WASP-39 b, HAT-P-1 b, and WASP-29 b with a SAD less than 0.990; HD 209458 b, WASP-74 b, HAT-P-26 b, and WASP-80 b with an SAD between 0.990 and 1.10; and the remaining eleven planets with an SAD greater than 1.110, among which GJ 436 b and GJ 3470 b return a SAD greater than 2. However, departing from this strict classification, we observe a more uniform distribution of SAD values across the Case 2 sample compared to the results of Case 1. With the exception of GJ 436 b and GJ 3470 b, which clearly exhibit a better fit with the stellar model, albeit with an extremely large error bar, the remaining spectra yield SAD values between 0.9 and 1.4. An SAD of 1.4 moderately favours the stellar model, but it is not significant enough to entirely disregard the base model. Given the large uncertainties associated with SAD values, we recommend caution before completely rejecting one model in favour of the other. We recognise that uncertainties in the SAD values typically range from 30-40% of the measurement value, occasionally exceeding the value itself, as observed in GJ 3470 b Case 2. The large SAD uncertainties stem from the considerable error bars of the STIS data, which, coupled with an inadequate overall fit resulting in a large 1 uncertainty in the model, contribute to high SAD errors. Particularly in the case of GJ 3470 b, despite the average data minus model distance being reduced by the stellar model, the significant uncertainties in both the STIS data and the retrieved stellar activity model lead to a considerable SAD error. Consequently, the error in the SAD measurement intrinsically implies that nearly all spectra for both spectral cases could potentially be best-fitted (or not) by a stellar contamination model. GJ 3470 b Case 3 and HD 209458 b Case 3 and 4 display a distance metric in the order of 1.0, indicating that both models effectively are able to fit the observed data in a comparable manner (Fig. 7).
@@ 5.4 SAT metric results
The Stellar Activity Temporal metric, or SAT in short, has been formulated to evaluate the consistency of observations, providing insights into the potential influence of stellar activity on spectral variations across different epochs. We report the SAT results in Fig. 8, where the transit depth has been converted into percentage points relative to the weighted average WFC3 data for each planet. The percentage average pairwise normalised differences among pairs of visits obtained with the same grating enables a direct comparison of the SATs for different planets, facilitating visual identification of those exhibiting the greatest spectral variations across epochs with respect to their atmospheric extents.
Similarly to the SAD, we observe a broad spectrum of SAT values, each representing distinct amplitudes of spectral temporal variations. Specifically, among the twenty-one SATs computed, eleven fall within the range of 1% to 4%. This indicates that over half of the planetary optical spectra exhibit largely consistent transit depth variations over time. Conversely, approximately seven out of twenty-one spectral combinations (relating to six planets since we identify here GJ 3470 b visit 1-visit 2 and GJ 3470 b visit 1-visit 3) yield a SAT between 6% and 12%, indicating a that another substantial subset of planets is characterised by more significant spectral variabilities. Within the lower-to-moderate variability regime, i.e. with an SAT falling within the 4-6% range, we identify WASP-6 b, GJ 3470 b visit 2-visit 3 (shown in Fig. 9), and WASP-74 b. It is important to mention that this categorisation is entirely arbitrary, derived from a visualisation of the SAT results in a histogram that reveals a bimodal distribution, roughly peaking at 2% and 7% (Fig. 10). HAT-P-26 b is the planet that displays the greatest variation among STIS visits, boasting a remarkable SAT variation of 11.1874.766%. By contrast, the planet with the smallest temporal variation is HD 209458 b G750L, featuring a SAT equal to 1.2380.296% (Fig. 11). While the two visits’ spectral features and overall modulations appear visually similar on HAT-P-26 b, especially redder than 0.45 m, the percentage scale reveals that the difference between data points at the same wavelength can reach up to 20%. Such discrepancy could potentially be significant for the retrieval to favour the stellar activity model over the base model. On the contrary, the G750L visits on HD 209458 b exhibit moderate disparities in spectral slopes between visits, yet the percentage discrepancy between any one pair of data points rarely exceeds 2% of the absolute transit depth.
@@
For certain planets such as WASP-17 b, GJ 3470 b and WASP-74 b a high SAT value corresponds to large variations in spectral modulations across visits, as the metric is naturally most sensitive to scenarios where the optical slopes display a significant change in gradient. Nevertheless, the rationale behind this metric is not solely to emphasise the distinct spectral shapes but rather to quantify the discrepancy between two visits and to flag temporal variations that warrant further monitoring and investigation, indicating potential activity within the system. Additionally, a low SAT value does not necessarily indicate low levels of stellar activity, as the star could be similarly active at subsequent epochs. Therefore, while the metric is not suitable for ruling out stellar activity, it can be used to inform follow-up studies. For instance, if a planet previously considered inactive exhibits a high SAT, we should not dismiss the possibility of activity altogether; rather, we should explore the system further.
@@ 6 Discussion
| Planet | Bayes Factor | SAD | SAT | ||||||
| Case 1 | Case 2 | Case 3 | Case 4 | Case 1 | Case 2 | Case 3 | Case 4 | G430L or other data combinations | |
| WASP-6 b | |||||||||
| WASP-52 b | |||||||||
| WASP-17 b | |||||||||
| WASP-121 b | |||||||||
| WASP-74 b | |||||||||
| HAT-P-26 b | |||||||||
| WASP-101 b | |||||||||
| HAT-P-11 b | |||||||||
| HAT-P-1 b | |||||||||
| WASP-31 b | |||||||||
| WASP-69 b | |||||||||
| WASP-80 b | |||||||||
| HAT-P-12 b | |||||||||
| WASP-29 b | |||||||||
| WASP-39 b | |||||||||
| HD 209458 b | (G430L); (G750L) | ||||||||
| HAT-P-18 b | |||||||||
| WASP-127 b | |||||||||
| GJ 3470 b | (V1-V3); (V1-V2); (V2-V3) | ||||||||
| GJ 436 b | |||||||||
@@
This study aims to characterise a variety of planets in a fully consistent way, both from a data analysis and atmospheric retrieval perspective. The comprehensive coverage provided by the combination of STIS and WFC3 data enabled us to explore the presence of various optical absorbers, cloud parameterizations, and the influence of stellar surface heterogeneities on planetary spectra. However, the quality and spectral coverage of the HST data available to us posed limitations on confidently identifying molecules beyond water. As the study progressed, it rapidly became evident that, at the population level, the spectra of these planets are predominantly affected by stellar activity. Stellar photospheric structures are capable of imprinting much stronger signatures on the planetary spectra compared to other atmospheric phenomena. Consequently, we redirected our focus towards assessing stellar contamination across the planetary sample, examining whether and to what extent the spectra of these planets are contaminated by the presence time-varying heterogeneities on the stellar photosphere. In Table 2 we present the collated stellar activity metric results for all of the planets in our sample. For the 15 planets whose observations are entirely consistent with one another we used standard competition-style ranking in an attempt to order them from most likely to be affected by stellar contamination to least likely. We stress that this is not a definitive assessment of stellar contamination; instead, it is merely a way of visually summarising the results of our study and highlighting the systems for which follow up, activity-oriented observations e.g. photometric monitoring surveys should be of highest priority. We also still include within Table 2 the 5 planets whose data sets are not consistent with the rest of the sample. However these are not included within the overall ranking as they are not directly comparable.
@@
Furthermore, in the sections below, we explore three of the planets in our population as in-depth case studies, putting the outcomes of our stellar activity metrics into the context of prior investigations, including supplementary stellar activity indicators, such as Ca H-K lines (S-indexes and values) and spot/facula crossing events identified in light curves. Our planet selection includes WASP-6 b, characterised by high levels of activity across all indicators; WASP-39 b, whose observations suggest a quiet stellar profile according to our activity metrics; and HAT-P-11 b, orbiting a notably active star, yet our findings consistently suggest that the host star may have been less active at the time of the STIS observations. The remaining planets included in our study undergo a more condensed analysis. We compile our findings comprehensively and present them in Appendix F for completeness.
@@ 6.1 WASP-6 b: Exploring a renowned active exoplanetary system
| H2O (base model) (sol 0) | - | - | - | 253.610.21 | |||
| H2O (base model) (sol 1) | 1.00260.0005 | - | - | - | 251.650.22 | ||
| H2O + clouds (sol 0) | - | - | 256.680.22 | ||||
| H2O + clouds (sol 1) | - | - | 253.070.22 | ||||
| H2O + clouds (sol 2) | - | - | 233.970.21 | ||||
| H2O + K + Na | - | 253.910.23 | |||||
| H2O + clouds + K + Na (sol 0) | 261.950.23 | ||||||
| H2O + clouds + K + Na (sol 1) | 255.410.24 | ||||||
| H2O + Stellar activity | - | - | - | 273.670.24 | |||
| H2O + Stellar activity + clouds | - | - | 273.430.24 | ||||
| H2O + Stellar activity + K + Na | - | 273.890.24 | |||||
| H2O + Stellar activity + clouds + K + Na | 273.180.25 | ||||||
@@
| H2O (base model) | - | - | - | 269.370.21 | |||
| H2O + clouds | - | - | 268.790.22 | ||||
| H2O + K + Na | - | 276.150.23 | |||||
| H2O + clouds + K + Na | 279.960.22 | ||||||
| H2O + Stellar activity | - | - | - | 274.940.23 | |||
| H2O + Stellar activity + clouds | - | - | 275.310.23 | ||||
| H2O + Stellar activity + K + Na | - | 278.230.24 | |||||
| H2O + Stellar activity + clouds + K + Na | 278.340.24 | ||||||
@@
On the other hand, the Case 2 spectrum comprising STIS G430L visit 2 displays a less steep optical slope. Due to the shallower slope and increased modulation in the optical data, a model incorporating clouds and alkali metals aligns most closely with the data exhibiting the highest evidence among all base model combinations. Both alkali metals abundances are well constrained and their distinct signatures are well identified in the spectrum by the retrieval. However, this model yields an isothermal temperature approximately 430 K lower than anticipated. In this context, introducing stellar activity to the water-only model shifts the retrieved temperature upwards, closer to the expected value given by the calculated equilibrium temperature. The addition of a layer of gray clouds to the base model and to the stellar activity model doesn’t significantly affect the retrieved parameters, as they remain consistent within a 1 range. From the posterior distributions of the stellar contamination model (Fig. 14), the retrieved cloud pressure of Pa suggests that, if present, the cloud deck tends towards higher pressures. Incorporating alkali metals enhances the model’s Bayesian evidence due to the presence of potential alkali features, yet their inclusion simultaneously results in a temperature decrease. It’s plausible that the model interprets the modulation within the data, coupled with the limited precision, as pronounced alkali features. Alternatively these broader modulations could feasibly also stem from the competing contamination effects of stellar spots and faculae if both were present on the photosphere. Discriminating between these two signatures with the current data quality poses a challenge. Ultimately, higher resolution data could aid in distinguishing between models more effectively.
@@
At first order the baseline stellar model improves the base model, as evidenced by the higher Bayesian evidence () associated with it. Compared to Case 1 and as shown in the posterior distribution (Fig. 14), the star in this scenario is expected to exhibit a slightly greater concentration of faculae ( = ) and fewer spots ( = ), as the former are necessary to slightly inhibit the strong positive bluewards optical slope imparted by spots. From the STIS observations, which are taken 6 days apart, the star appears to be consistently active over that timescale, albeit with varying degrees of activity. The activity level appearing to remain relatively high over this period would be consistent with the stellar rotation period of 23.800.15 days determined in Bonomo et al. (2017). Our analysis of the stellar activity indicators reveals that WASP-6 b exhibits a Bayes Factor of approximately 20 for Case 1 and above 5 for Case 2, strongly favouring the inclusion of stellar contamination in both cases. Similarly, with a SAD = 2.9721.652 for Case 1 and 1.2620.529 for Case 2, WASP-6 emerges as the most active star among all Case 1 scenarios and amongst the most active for Case 2. The SAT is also notably high at 5.919%1.494%, emphasising the diversities amongst the two G430L visits, which likely requires a degree of active region evolution or decay to fully explain. A visual inspection of the STIS white light curves doesn’t definitively reveal any instances of stellar spot or faculae crossings, although it’s challenging to ascertain their presence due to gaps in the light curves. The active nature of the star has been confirmed extensively in the literature. Pinhas et al. (2018) mention that WASP-6 b is the planet with the highest level of stellar activity contamination among their whole sample. High-resolution spectra of the stellar host obtained with HARPS, revealed indications of chromospheric activity. Within the wavelength region encompassing the Caii H&K lines, all HARPS spectra exhibit evidence of emission, suggesting a moderate level of stellar activity (Nikolov et al., 2015). Interestingly, despite the presence of such activity, the three Hubble Space Telescope visits analysed by Nikolov et al. (2015) demonstrated consistent stellar flux levels, and the corresponding light curves displayed no discernible evidence of spot-crossing events. Furthermore, literature values of corroborate our suggestion that WASP-6 is moderate-to-highly active, Sing et al. (2016) report a value of = -4.741 while Carter et al. (2020) find a value of = -4.5110.037, which was derived from the HARPS spectra presented in Gillon et al. (2009). In summary, WASP-6 appears to be a star with moderate activity when compared to other stars of similar spectral type (Boro Saikia et al., 2018; Carter et al., 2020).

### v3

@@ 5.3 SAD metric results
As discussed in Section 5.2, the Bayesian approach favours precise data, giving less weight to data points with larger uncertainties. Attempting a retrieval on the spectral bands covered by STIS data only is not advisable, as the inclusion of WFC3 data is necessary to anchor the spectra to the continuum level. However, we still want to assess how well the retrieval captures potential stellar activity signals in the STIS data sets. Hence, we utilise the SAD metric to assesses the goodness-of-fit in the optical wavelengths, which acts independently of Bayesian statistics. A description of the metric is given in Section 4.2 while the results are reported in Fig. 5 and 6 for Case 1 and Case 2 respectively, in ascending SAD order. The SAD results for GJ 3470 b Case 3 and HD 209458 b Case 3 and 4 are displayed in Fig. 7.
Based on the SAD Case 1 results, it can be inferred that WASP-127 b is the planet for which the stellar contamination model provides the least accurate fit to the optical data, likely due to the significant scatter in the data points. Conversely, the strong optical slope observed in WASP-6 b’s STIS data is much better explained by the activity model than by the base model. Between WASP-127 b and WASP-6 b, a wide range of SAD values reflects the diversity in spectral shapes and models’ ability to fit the data. Six planets (HD 209458 b, WASP-69 b, HAT-P-11 b, HAT-P-12 b, WASP-31 b, and WASP-29 b) have SAD values near 1 (0.99SAD1.10), indicating negligible/weak preference for either model. WASP-127 b and WASP-39 b mostly favour the base model (SAD 0.99), while planets like GJ 3470 b, WASP-80 b, HAT-P-26 b and others are better described by an active star model (SAD 1.10).
In Case 2, applying similar criteria, WASP-39 b, HAT-P-1 b, and WASP-29 b show SAD values below 0.99, while HD 209458 b, WASP-74 b, HAT-P-26 b, and WASP-8 b fall within the 0.99 to 1.10 range. The remaining eleven planets exceed a 1.10 value for the SAD, with GJ 436 b and GJ 3470 b notably higher than 2. Case 2 exhibits a more uniform SAD distribution compared to Case 1. Except for GJ 436 b and GJ 3470 b, which clearly prefer the stellar model but with large error bars, the rest of the spectra yield SAD values between 0.9 and 1.4. SAD values around 1.4 moderately favour the stellar model, but not decisively. For GJ 3470 b Case 3 and HD 209458 b Cases 3 and 4, displayed in Fig. 7, SAD values around 1.0 suggest both models fit the data comparably well.
The considerable uncertainties in SAD values, often 30-40% or more, warrant caution when favouring one model over the other. High SAD errors result from large STIS data error bars coupled with poor model fits. This is particularly evident in GJ 3470 b Case 2, where even though the stellar model reduces the average data-model distance, substantial uncertainties in both parameters lead to a considerable SAD error. Given these large uncertainties, it is possible that any spectrum in the sample could be best fitted by either model.
@@ 5.4 SAT metric results
The Stellar Activity Temporal metric, or SAT in short, has been formulated to evaluate the consistency of observations, providing insights into the potential influence of stellar activity on spectral variations across different epochs. We report the SAT results in Fig. 8, where the transit depth has been converted into percentage points relative to the weighted average WFC3 data for each planet. The percentage SAT enables direct comparisons among planets, facilitating visual identification of those exhibiting the greatest spectral variations across epochs with respect to their atmospheric extents.
Similarly to the SAD, we observe a broad spectrum of SAT values, each representing distinct amplitudes of spectral temporal variations. Specifically, among the twenty-one SATs computed, eleven fall within the range of 1% to 4%. This indicates that over half of the planetary optical spectra exhibit largely consistent transit depth variations over time. On the other hand, about seven spectral combinations (involving six planets, as these include GJ 3470 b visit 1-visit 2 and GJ 3470 b visit 1-visit 3) out of twenty-one show a SAT between 6% and 12%. This indicates that another substantial subset of planets is characterised by more significant spectral variabilities. Within the lower-to-moderate variability regime, i.e. with an SAT falling within the 4-6% range, we identify WASP-6 b, GJ 3470 b visit 2-visit 3 (shown in Fig. 9), and WASP-74 b. It is important to mention that this categorisation is entirely arbitrary, derived from a visualisation of the SAT results in a histogram that reveals a bimodal distribution, roughly peaking at 2% and 7% (Fig. 10). HAT-P-26 b is the planet that displays the greatest variation among STIS visits, boasting a remarkable SAT variation of 115%. By contrast, the planet with the smallest temporal variation is HD 209458 b G750L, featuring a SAT equal to 1.20.3% (Fig. 11). While the two visits’ spectral features and overall modulations appear visually similar on HAT-P-26 b, especially redder than 0.45 m, the percentage scale reveals that the difference between data points at the same wavelength can reach up to 20%. Such discrepancy could potentially be significant for the retrieval to favour the stellar activity model over the base model. On the contrary, the G750L visits on HD 209458 b exhibit moderate disparities in spectral slopes between visits, yet the percentage discrepancy between any one pair of data points rarely exceeds 2% of the absolute transit depth.
@@
For certain planets such as WASP-17 b, GJ 3470 b and WASP-74 b a high SAT value corresponds to large variations in spectral modulations across visits, as the metric is naturally most sensitive to scenarios where the optical slopes display a significant change in gradient. Nevertheless, the rationale behind this metric is not solely to emphasise the distinct spectral shapes. It is also intended to quantify the discrepancy between two visits and to flag temporal variations that warrant further monitoring and investigation, indicating potential activity within the system. Additionally, a low SAT value does not necessarily indicate low levels of stellar activity, as the star could be similarly active at subsequent epochs. Therefore, while the metric is not suitable for ruling out stellar activity, it can be used to inform follow-up studies. For instance, if a planet previously considered inactive exhibits a high SAT, we should not dismiss the possibility of activity altogether; rather, we should explore the system further.
@@ 6 Discussion
This study aims to characterise a variety of planets in a fully consistent way, both from a data analysis and atmospheric retrieval perspective. The comprehensive coverage provided by the combination of STIS and WFC3 data enabled us to explore the presence of various optical absorbers, cloud parameterisations, and the influence of stellar surface heterogeneities on planetary spectra. However, the quality and spectral coverage of the HST data available to us posed limitations on confidently identifying molecules beyond water. As the study progressed, it rapidly became evident that, at the population level, the spectra of these planets are predominantly affected by stellar activity. Stellar photospheric structures are capable of imprinting much stronger signatures on the planetary spectra compared to other atmospheric phenomena. Consequently, we redirected our focus towards assessing stellar contamination across the planetary sample, examining whether and to what extent the spectra of these planets are contaminated by the presence time-varying heterogeneities on the stellar photosphere. In Fig. 12 we present the collated stellar activity metric results for all of the planets in our sample. For the 15 planets whose observations are entirely consistent with one another we used standard competition-style ranking in an attempt to order them from most likely to be affected by stellar contamination to least likely. We stress that this is not a definitive assessment of stellar contamination; instead, it is merely a way of visually summarising the results of our study and highlighting the systems for which follow up, activity-oriented observations e.g. photometric monitoring surveys should be of highest priority. We also still include within Fig. 12 the 5 planets whose data sets are not consistent with the rest of the sample. However these are not included within the overall ranking as they are not directly comparable.
@@
Furthermore, in the sections below, we explore three of the planets in our population as in-depth case studies, putting the outcomes of our stellar activity metrics into the context of prior investigations, including supplementary stellar activity indicators, such as Ca H&K lines (S-indexes and values) and spot/facula crossing events identified in light curves. Our planet selection includes WASP-6 b, characterised by high levels of activity across all indicators; WASP-39 b, whose observations suggest a quiet stellar profile according to our activity metrics; and HAT-P-11 b, orbiting a notably active star, yet our findings consistently suggest that the host star may have been less active at the time of the STIS observations. The remaining planets included in our study undergo a more condensed analysis. We compile our findings comprehensively and present them in Appendix F for completeness.
@@ 6.1 WASP-6 b: exploring a renowned active exoplanetary system
H2O (base model) (sol 0) | - | - | - | 253.60.2 | |||
H2O (base model) (sol 1) | - | - | - | 251.70.2 | |||
H2O + clouds (sol 0) | - | - | 256.70.2 | ||||
H2O + clouds (sol 1) | - | - | 253.10.2 | ||||
H2O + clouds (sol 2) | - | - | 234.00.2 | ||||
H2O + K + Na | - | 253.90.2 | |||||
H2O + clouds + K + Na (sol 0) | 262.00.2 | ||||||
H2O + clouds + K + Na (sol 1) | 255.40.2 | ||||||
H2O + Stellar activity | - | - | - | 273.70.2 | |||
H2O + Stellar activity + clouds | - | - | 273.40.2 | ||||
H2O + Stellar activity + K + Na | - | 273.90.2 | |||||
H2O + Stellar activity + clouds + K + Na | 273.20.3 | ||||||
@@
H2O (base model) | - | - | - | 269.40.2 | |||
H2O + clouds | - | - | 268.80.2 | ||||
H2O + K + Na | - | 276.20.2 | |||||
H2O + clouds + K + Na | 280.00.2 | ||||||
H2O + Stellar activity | - | - | - | 274.90.2 | |||
H2O + Stellar activity + clouds | - | - | 275.30.2 | ||||
H2O + Stellar activity + K + Na | - | 278.20.2 | |||||
H2O + Stellar activity + clouds + K + Na | 278.30.2 |
@@
On the other hand, the Case 2 spectrum comprising STIS G430L visit 2 displays a less steep optical slope. Due to the shallower slope and increased modulation in the optical data, a model incorporating clouds and alkali metals aligns most closely with the data by exhibiting the highest evidence among all base model combinations. Both alkali metals abundances are well constrained and their distinct signatures are well identified in the spectrum by the retrieval. However, this model yields an isothermal temperature approximately 430 K lower than anticipated. In this context, introducing stellar activity to the water-only model shifts the retrieved temperature upwards, closer to the expected value given by the calculated equilibrium temperature. The addition of a layer of gray clouds to the base model and to the stellar activity model doesn’t significantly affect the retrieved parameters, as they remain consistent within a 1 range. From the posterior distributions of the stellar contamination model (Fig. 15), the retrieved cloud pressure of Pa suggests that, if present, the cloud deck tends towards higher pressures. Incorporating alkali metals enhances the model’s Bayesian evidence due to the presence of potential alkali features, yet their inclusion simultaneously results in a temperature decrease. It is plausible that the model interprets the modulation within the data, coupled with the limited precision, as pronounced alkali features. Alternatively these broader modulations could feasibly also stem from the competing contamination effects of stellar spots and faculae if both were present on the photosphere. Discriminating between these two signatures with the current data quality poses a challenge. Ultimately, higher resolution data could aid in distinguishing between models more effectively.
@@
At first order the baseline stellar model improves the base model, as evidenced by the higher Bayesian evidence () associated with it. Compared to Case 1 and as shown in the posterior distribution in Fig. 15, the star in this scenario is expected to exhibit a slightly greater concentration of faculae ( = ) and fewer spots ( = ). The former are necessary to slightly inhibit the strong positive bluewards optical slope imparted by spots. From the STIS observations, which were taken 6 days apart, the star appears to have been consistently active over that timescale, albeit with varying degrees of activity. The activity level appearing to remain relatively high over this period would be consistent with the stellar rotation period of 23.80.2 days determined in Bonomo et al. (2017). Our analysis of the stellar activity indicators reveals that WASP-6 b exhibits a Bayes factor of approximately 20 for Case 1 and above 5 for Case 2, strongly favouring the inclusion of stellar contamination in both cases. Similarly, with a SAD = 32 for Case 1 and 1.30.5 for Case 2, WASP-6 emerges as the most active star among all Case 1 scenarios and amongst the most active for Case 2. The SAT is also notably high at 6%1%, emphasising the diversities amongst the two G430L visits, which likely requires a degree of active region evolution or decay to fully explain. A visual inspection of the STIS white light curves doesn’t definitively reveal any instances of stellar spot or faculae crossings, although it’s challenging to ascertain their presence due to gaps in the light curves. The active nature of the star has been confirmed extensively in the literature. Pinhas et al. (2018) mention that WASP-6 b is the planet with the highest level of stellar activity contamination among their whole sample. High-resolution spectra of the stellar host obtained with HARPS, revealed indications of chromospheric activity. Within the wavelength region encompassing the Caii H&K lines, all HARPS spectra exhibit evidence of emission, suggesting a moderate level of stellar activity (Nikolov et al., 2015). Interestingly, despite the presence of such activity, the three Hubble Space Telescope visits analysed by Nikolov et al. (2015) demonstrated consistent stellar flux levels, and the corresponding light curves displayed no discernible evidence of spot-crossing events. Furthermore, literature values of corroborate our suggestion that WASP-6 is moderate-to-highly active: Sing et al. (2016) report a value of = -4.741 while Carter et al. (2020) find a value of = -4.510.04, which was derived from the HARPS spectra presented in Gillon et al. (2009). In summary, WASP-6 appears to be a star with moderate activity when compared to other stars of similar spectral type (Boro Saikia et al., 2018; Carter et al., 2020).